OpenStudy (anonymous):

Given that f(x) = x^^2 − 4x − 3 and g(x) = the quantity of x plus three, over four , solve for f(g(x)) when x = 9. −6 3 6 9

OpenStudy (anonymous):

any ideas

OpenStudy (anonymous):

it would help if you put the steps js

OpenStudy (anonymous):

g(x) = the quantity of x plus three, over four , First try tol translate this into maths. When you are done, find g(x) when x=9 You will get a numerical solution (something like g(x)=1) Put this solution into: f(x) = x^2 − 4x − 3 You are done!

OpenStudy (anonymous):

the equation is f(x) = x^2 − 4x − 3

OpenStudy (anonymous):

can you show me in steps

OpenStudy (anonymous):

The steps are there, try to work it through

OpenStudy (anonymous):

f(9) = (9)^2 - 4(9) - 3 f(9) = 81 -36 -3 f(9) =42 is this correct

OpenStudy (anonymous):

i dont get how to do the g(x)

OpenStudy (anonymous):

Do NOT jump steps. Translate that English sentence into maths first!

OpenStudy (anonymous):

g(x) = x+3/4

OpenStudy (anonymous):

(x+3)/4

OpenStudy (anonymous):

g(x) = x+3/4\[g(x) = \]

OpenStudy (anonymous):

So what will be g(x) at x=9?

OpenStudy (anonymous):

so i subsitute the 9 for the x

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

3

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

So that is g(x) at x=9 What will be f(g(x)) at x=9?

OpenStudy (anonymous):

the answer isnt 3

OpenStudy (anonymous):

No :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3 is g(x) at x=9

OpenStudy (anonymous):

do i multiply the answer to f() and G(x)

OpenStudy (anonymous):

f(g(x)) at x=9 is the question! G(x)=3 So f(g(x)) at x=9? is f(3)

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

there are 2 functions f(x) and g(x) f(g(x)) at x=9 is the question f(g(x)) means you have x=9 and you first do function g than you get some result (3 in our case) and you carry on to do function f with your result. f(g(x))=f(g(9))=f(3)=3^2-4*3-3=-6

OpenStudy (anonymous):

the answer is -6 then

OpenStudy (anonymous):

It seems you do not wish to understand just looking for an answer... Yes it is -6